OpenStudy (anonymous):

Factor the polynomial 16x^2-48x+36

OpenStudy (anonymous):

oh wait...i did that wrong..

OpenStudy (anonymous):

This is a perfect square trinomial. To find the factors, first look at the first and last terms. Notice that 16x^2 is (4x)^2 and 36 is 6^2. The middle term is the result of taking twice the product of these two terms, resulting in a perfect square trinomial. It factors as (4x+6)^2 Which can further be factored into (2(2x+3))^2 Finally, this simplifies to 4(2x+3)^2.

OpenStudy (zehanz):

If it was of the form a²-2ab+b², it could be written as (a-b)². It certainly LOOKS like this! After all, 16x² is the square of 4x, 36 is 6², so it could well be (4x-6)². If we check the middle term (which should be -2ab) we have -2*4x*6 = -48x, so it is just right! Now one extra number can be factored out: (4x-6)²=(2(2x-3))²=4(x-3)² So my final answer would be: 4(x-3)²

OpenStudy (anonymous):

Forgot the coefficient of x is 2, ZeHanz.

OpenStudy (zehanz):

Right you are, @calmat01 ! 4(2x-3)²

OpenStudy (anonymous):

so it would be \[4(2x-3)^2\]

OpenStudy (anonymous):

yes. That is correct!

OpenStudy (anonymous):

how would I show my work on this one?

OpenStudy (anonymous):

I got the answer, but it would be nice if I had some work shown to backer up haha!

OpenStudy (anonymous):

Never mind I got it! Thank you so much, this website has really been helping me out!